8.2 What is/are the critical value(s) of the z-test statistic for this two-tailed hypothesis test, where α=0.05? Use the appropriate value(s) from the table. Use a comma and a space to separate answers as needed.

z0.10 z0.05 z0.025 z0.01 z0.005
1.282 1.645 1.960 2.326 2.576

Critical Value (s) :
Answer #1

8.2)

Solution :

  = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z 0.025 = 1.96

Critical values are : -1.96 and +1.96